Find the eigenvalues and eigenfunctions of the given boundary value problem + Ag = 0, / (0) = 0, 4( L) = 0Please clearly show all work. Thank you.
Answer #1

Given boundary Value problem is Cane I-o then forubion. of * ia ato=O C20 0 6tte trivial souion of x Canel = -H(t0. Tan * beae L Tf A70 Put = 2 (p+. Then * belomes Icharacteristc puation is +H0 GCa Kt) ut) b D)0> C2=0 C2 = 0 n(Ht2lkt) Har nemivial S